# Imaging the indirect-to-direct band-gap crossover in PbI2

### v1

@@ Abstract
The nature of the band-gap in PbI2 (i.e. direct or indirect) is crucial for its applications. Here we directly image, using angle-resolved photoemission spectroscopy, its thickness-dependent crossover from an indirect to a direct band-gap. We experimentally probe a large shift of the valence band maximum towards the center of the Brillouin zone, when the thickness of PbI2 films is greater than a monolayer. Our experimental results, accompanied by density functional theory calculations, suggest that the band-gap crossover is driven by interlayer interactions and the hybridization of iodine $p_{z}$ orbitals. These findings demonstrate the tunable electronic properties of PbI2, highlighting its potential for applications in optoelectronics.
@@
The nature of the band-gap in a given semiconductor is crucial for its applications: a photon with an energy of the order of 1 eV, close to most of semiconductors band-gaps, carries a momentum of the order of $5\times 10^{-4}$-1. This value is negligible compared to the typical separation, in reciprocal space, between the valence band maximum (VBM) and the conduction band minimum (CBM) in indirect-gap semiconductors. Thus, a direct band-gap favors optical absorption and radiative emission (ideal for lasers and LEDs [1, 2, 3, 4]), but also results in shorter electron-hole pair lifetimes. On the other hand, an indirect band-gap, where optical transitions between the valence and conduction bands require the assistance of phonons, results in less efficient optical absorption (increased transparency), but also yields an increase in exciton lifetimes through the inhibition of spontaneous radiative emission. Increased carrier lifetimes and the absence of spontaneous light emission open up applications in high-efficiency solar cells [5, 6] and low-optical-noise photonic circuits [7, 8].
@@
Two-dimensional layered semiconductors have attracted significant research interest as platforms for studying fundamental quantum phenomena [9, 10, 11, 12, 13], but most importantly due to their applications in nanoelectronics and optoelectronics [14, 15, 16, 17, 18, 19, 20, 21]; an application potential that is a closely related to the nature of their band-gap. Among the various 2D layered semiconductors, lead iodide (PbI2) stands out with wide use in X-ray detectors [22, 23, 24], and as a precursor for organolead halide perovskites that offer considerable prospects in light-emitting diodes [25, 26], photodetectors [27, 28] and solar cells [29, 30]. A peculiar characteristic of PbI2 is that its direct band-gap, the key feature behind its applications, becomes indirect when its thickness decreases to a single monolayer [31, 32, 33, 34, 35, 36]. This band-gap crossover in PbI2 strongly affects its optoelectronic and vibrational properties [31, 32, 33], while opening up a different spectrum of applications [37, 33, 38, 39]. The thickness-dependent band-gap crossover in 2D layered semiconductors is, however, not exclusive to PbI2. Transition metal dichalcogenides (TMDs), also, exhibit a change in the nature of their band-gaps in the monolayer form [40, 41, 42, 43, 44]. The band-gap in TMDs follows, however, an opposite trend than in PbI2, with a direct band-gap in monolayer TMD films and an indirect gap in their bulk form.
Figure 1: Structural and electronic properties of the 2H polytype of PbI2. (a) Atomic structure of PbI2, showing a sandwich-like arrangement where a layer of Pb atoms (grey) is positioned between two layers of I atoms (purple and pink). (b) Top view of the monolayer atomic structure, revealing a hexagonal arrangement of iodine atoms surrounding each lead atom. (c) Bulk and surface Brillouin zone with the $k_{z}=0$ and $k_{z}=\pi/c$ planes highlighted in green and blue, respectively. (d–g) Theoretical band structure of PbI2 obtained from GGA PBE calculations, for different number of layers. The calculations were performed for free-standing PbI2. For multilayers, a slab approach has been used, thus resulting into a multitude of bands that evolve into a continuum in thicker films. Red arrows indicate the position of the minimum band-gap, with tilted (vertical) arrows denoting indirect (direct) band-gaps. The energy axis of each panel has been scaled with respect to its valence band maximum.
@@
The thickness-dependent band-gap crossover in PbI2 has been theoretically predicted by band structure calculations, which agree that for a monolayer film the VBM is located away from the center of the Brillouin zone, while it shifts to the latter when interlayer coupling comes into play, thereby altering the electronic structure in multilayer films [31, 32, 33, 34, 35, 36]. The mechanism for the band-gap crossover, should therefore have a direct fingerprint in the experimental electronic structure of PbI2. However, to this day, the experimental verification of the band-gap crossover only comes by indirect experimental means, such as photoluminescence and optical measurements that lack sensitivity in momentum space [33, 36, 45].
@@
Here, we present direct experimental evidence on the electronic structure of PbI2 for the indirect-to-direct band-gap crossover as its thickness increases beyond a single monolayer. By means of angle-resolved photoemission spectroscopy (ARPES), we investigated the electronic structure of PbI2 thin films deposited on graphene/6H-SiC(0001) [46, 47]. Through a systematic study of films with different thicknesses, we were able to track the evolution of the experimental electronic structure and reveal the change in the nature of the band-gap. Our experimental results are in excellent agreement with density functional theory (DFT) calculations and pinpoint the exact location of the energy gap in reciprocal space. Detailed information on the experimental technique, sample growth, and DFT calculations is provided in section SM1 of the Supplemental Material  [48].
@@
The crystal structure of PbI2 is schematically depicted in Figs. 1(a) and 1(b). The PbI2 monolayer adopts a hexagonal in-plane arrangement with Pb atoms sandwiched between two layers of I atoms. Each Pb atom is surrounded by six I atoms, forming a nearly octahedral environment. The bulk structure consists of stacked layers held together by weak van der Waals forces, forming the 2H polytype, the most stable phase of PbI2 [49, 50, 51, 52, 53, 54]. Bulk PbI2 crystallizes in the $P\bar{3}m1$ space group (No. 164) with lattice parameters $a=b=4.56$ and $c=6.99$ [55]. The surface and bulk Brillouin zones (BZs) are presented in Fig. 1(c), along with their high-symmetry points. In the following, we will discuss the electronic structure of PbI2 in films of various thicknesses.
Figure 2: Electronic structure of 1 layer (1L) and 7 layers (7L) PbI2 films. (a) and (b) Band structure of, respectively, 1L and 7L films along the $\bar{\text{K}}$-$\bar{\Gamma}$-$\bar{\text{M}}$-$\bar{\text{K}}$ path. (c) and (d) Band structure of, respectively, 1L and 7L films with maximally enhanced contrast for each momentum distribution curve (MDC). The contrast enhancement was achieved by normalizing each MDC to its integrated intensity. The tail of the red arrows indicates the location of the VBM for both films. (e) and (f) Constant energy maps corresponding to the binding energy of $1.8$ eV (i.e. at the valence band maximum) for 1L and 7L films, respectively, under identical measurement conditions. (g) Energy distribution curves (EDCs) corresponding to panels (c) and (d), obtained by integrating over a momentum window of 0.1 Å-1 around $\bar{\Gamma}$. The blue (black) curve corresponds to 7L (1L) films. Measurements have been performed at $16$K using photons of $80$ eV and linear horizontal polarization.
@@
The most important change in the electronic structure as a function of film thickness is the emergence of a direct band-gap, which has an immediate impact on the properties and technological applications of PbI2 [56, 57, 58]. Before discussing the experimental results, Fig. 1(d)-(f) shows the evolution of the electronic structure with an increasing number of PbI2 layers, as predicted by DFT calculations. For a single layer [panel (a)] the band-gap is indirect due to the fact that the VBM is not located at a high-symmetry point of the surface Brillouin zone (SBZ), but rather in-between $\bar{\Gamma}$ and $\bar{\text{M}}$. As the thickness increases [Figs. 1(e) and 1(g)], the VBM shifts towards $\bar{\Gamma}$ giving rise to a direct band-gap for films of three layers (3L) and beyond. The calculated gap size slightly varies as a function of thickness, ranging from $\sim$1.95 eV for a single layer, to $\sim$1.54 eV for 7L films. Since the location of the CBM is not affected by the increasing number of layers, we underline the key role of the VBM in establishing the nature of the band-gap: as the thickness increases beyond 1L, the emergence of electronic bands with a maximum at $\bar{\Gamma}$ determines the position of the VBM, which stays at $\bar{\Gamma}$ even after the development of the bulk continuum.
@@
Previous theoretical work has demonstrated that the top of the valence band in single-layer PbI2 stems from in-plane I $5p$ orbitals [32]. However, as the thickness increases, interactions between neighboring layers result into the emergence of the aforementioned electronic bands, of $p_{z}$ character, that exhibit a maximum at $\bar{\Gamma}$. The shift of the VBM to $\bar{\Gamma}$ is obvious when the film thickness is at least 3L [panels (f) and (g)], with the bilayer film [panel (e)] marking an intermediate stage between the monolayer behavior and the electronic structure of thicker multilayer films. The indirect-to-direct band-gap crossover of PbI2 is in agreement with previous theoretical studies [32, 36, 33, 34]. It has also been used to explain the findings of experimental measurements with no direct view on the electronic structure of PbI2  [33, 45]. A direct experimental verification of the indirect-to-direct band-gap crossover has never been performed and it will be the focus of the next paragraphs.

### v2

# Dimensional control of the band-gap crossover in layered lead iodide

@@ Abstract
Before assessing the suitability of a semiconductor for specific applications, the first question to ask is whether it possesses a direct or indirect band-gap. This distinction is fundamental, as the operation of devices such as light-emitting diodes, solar cells, and photodetectors is closely tied to the band-gap nature. Semiconductors that exhibit a band-gap crossover - indirect to direct or vice versa - offer enhanced versatility for optoelectronic applications. Prominent examples include transition metal dichalcogenides and the subject of this study, PbI2. The nature of the band-gap, and its crossover, can only be directly determined in reciprocal space by tracking the valence-band maximum and conduction-band minimum. Here, we directly visualize the thickness-dependent crossover of PbI2 from an indirect to a direct band-gap using angle-resolved photoemission spectroscopy. Our measurements reveal a shift of the valence-band maximum toward the Brillouin-zone center as the film thickness exceeds a monolayer. Supported by density functional theory calculations, our results show that this crossover is driven by interlayer interactions and the hybridization of iodine $p_{z}$ orbitals. These findings demonstrate the tunable electronic structure of PbI2 and its potential for optoelectronic applications.
I Introduction
@@
The nature of the band-gap in a given semiconductor is crucial for its applications: a photon with an energy of the order of 1 eV, close to band gaps of most semiconductors, carries a momentum of the order of $5\times 10^{-4}$ Å-1. This value is negligible compared to the typical Brillouin-zone sizes of solids (of the order of $1$ Å-1) and to the separation, in reciprocal space, between the valence band maximum (VBM) and the conduction band minimum (CBM) in indirect-gap semiconductors. Thus, a direct band-gap favors optical absorption and radiative emission (ideal for lasers and LEDs [50, 57, 24, 14]), but also results in shorter electron-hole pair lifetimes. On the other hand, an indirect band-gap, where optical transitions between the valence and conduction bands require the assistance of phonons, results in less efficient optical absorption (increased transparency), but also yields an increase in exciton lifetimes through the inhibition of spontaneous radiative emission. Increased carrier lifetimes and the absence of spontaneous light emission open up applications in Si-based solar cells [38, 73], mechanical sensors [60, 49] and low-optical-noise photonic circuits [42, 1].
@@
Two-dimensional layered semiconductors have attracted significant research interest as platforms for studying fundamental quantum phenomena [10, 39, 76, 8, 51], but most importantly due to their applications in nanoelectronics and optoelectronics [15, 79, 75, 12, 17, 58, 11, 22]; an application potential that is closely related to the nature of their band-gap. Among the various 2D layered semiconductors, lead iodide (PbI2) stands out with wide use in X-ray detectors [61, 62, 43], and as a precursor for organolead halide perovskites that offer considerable prospects in light-emitting diodes [70, 21], photodetectors [67, 25] and solar cells [35, 18]. A peculiar characteristic of PbI2 is that its direct band-gap, the key feature behind its applications, becomes indirect when its thickness decreases to a single monolayer [63, 80, 82, 77, 74, 71]. This band-gap crossover in PbI2 strongly affects its optoelectronic and vibrational properties [63, 80, 82], while opening up a different spectrum of applications [23, 82, 34, 72]. The thickness-dependent band-gap crossover in 2D layered semiconductors is, however, not exclusive to PbI2. Transition metal dichalcogenides (TMDs), also, exhibit a change in the nature of their band-gaps in the monolayer form [26, 40, 68, 41, 81]. The band-gap in TMDs follows, however, an opposite trend than in PbI2, with a direct band-gap in monolayer TMD films and an indirect gap in their bulk form.
@@
The thickness-dependent band-gap crossover in PbI2 has been theoretically predicted by band structure calculations, which agree that for a monolayer film the VBM is located away from the center of the Brillouin zone, while it shifts to the latter when interlayer coupling comes into play, thereby altering the electronic structure in multilayer films [63, 80, 82, 77, 74, 71]. The mechanism for the band-gap crossover should therefore have a direct fingerprint in the experimental electronic structure of PbI2. However, to this day, the experimental verification of the band-gap crossover only comes by indirect experimental means, such as photoluminescence and optical measurements that lack sensitivity in momentum space [82, 71, 66].
@@
Here, we present direct experimental evidence on the electronic structure of PbI2 for the indirect-to-direct band-gap crossover as its thickness increases beyond a single monolayer. By means of angle-resolved photoemission spectroscopy (ARPES), we investigated the electronic structure of PbI2 thin films deposited on graphene/6H-SiC(0001) [54, 5]. Through a systematic study of films with different thicknesses, we were able to track the evolution of the experimental electronic structure and reveal the change in the nature of the band-gap. Our experimental results are in excellent agreement with density functional theory (DFT) calculations and pinpoint the exact location of the energy gap in reciprocal space.
II Methods
II.1 Sample growth
Figure 1: Growth dynamics of 12 layers of PbI2 on graphene/SiC(0001), probed by a 600 eV He beam aligned along the graphene zigzag direction. (a) Evolution of the GIFAD pattern. (b) Reflectivity curve.
The sample growth was conducted in a homemade Molecular Beam Epitaxy chamber at a base pressure in the low $10^{-10}$ mbar range. Graphene substrates, grown on 6H-SiC(0001), were prepared by annealing cycles at $520\,^{\circ}$C until the observation of a well-contrasting GIFAD pattern. The deposition was carried out by heating the PbI2 crucible at $300\,^{\circ}$C, while the substrate is at room temperature.
The growth dynamics was monitored in real time by Grazing Incidence Fast Atom Diffraction (GIFAD) [29, 3]. GIFAD relies on the coherent scattering of keV-energy He atoms at incidence angles typically lower than $1^{\circ}$ with respect to the surface plane. In contrast to X-ray or electron diffraction, GIFAD is a very soft technique that allows sensitive and damage-free characterization of the topmost layer only, with unambiguous information on film thickness. The latter is directly accessible through the intensity oscillations of the reflected atoms. Fig. 1(a) shows the time evolution of the GIFAD pattern during the growth of a 12-layer thick PbI2 sample. The probe beam was parallel to the zigzag direction of graphene. The well-resolved structure along $\Delta k$ demonstrates a good relative alignment between the PbI2 and substrate lattices [66], with armchair (PbI2) $\parallel$ zigzag (graphene). The sharp contrast of the time-dependent oscillations, reproduced in Fig. 1(b), indicates a layer-by-layer growth mode, while the high intensity, relative to that measured on the bare substrate, reflects the high quality of the sample.
II.2 Angle-resolved photoemission spectroscopy
Angle-resolved photoemission measurements were performed at the CASSIOPEE beamline of Synchrotron SOLEIL, France. The samples were transported in a UHV suitcase, directly after growth, with the pressure never exceeding $5\times 10^{-8}$ mbar during the transfer. During the ARPES measurements, the pressure remained below $3\times 10^{-10}$ mbar, and the sample temperature was stabilized at 16 K. Data were acquired using photon energies of 80 eV and 65 eV and linear horizontal polarization. The energy resolution was maintained at 30 meV.
Figure 2: Structural and electronic properties of the 2H polytype of PbI2. (a) Atomic structure of PbI2, showing a sandwich-like arrangement where a layer of Pb atoms (grey) is positioned between two layers of I atoms (purple and pink). (b) Top view of the monolayer atomic structure, revealing a hexagonal arrangement of iodine atoms surrounding each lead atom. (c) Bulk and surface Brillouin zone with the $k_{z}=0$ and $k_{z}=\pi/c$ planes highlighted in green and blue, respectively. (d)–(g) Theoretical band structure of PbI2 obtained from GGA PBE calculations, for different number of layers. The calculations were performed for free-standing PbI2. For multilayers, a slab approach has been used, thus resulting in a multitude of bands that evolve into a continuum in thicker films. Red arrows indicate the position of the minimum band-gap, with tilted (vertical) arrows denoting indirect (direct) band-gaps. The energy axis of each panel has been scaled with respect to its valence band maximum.
II.3 First-principles calculations
First principles density functional theory (DFT) calculations were performed using the projected augmented-wave (PAW) [6] potentials implemented in the Vienna Ab initio Simulation Package (VASP) [31, 30, 32] code. For the exchange-correlation energy, the generalized gradient approximation (GGA) in the Perdew–Burke–Ernzerhof (PBE) parametrization [56] was used. The energy cutoff for the plane-wave expansion was set to $400$ eV. Furthermore, the van der Waals correction was introduced within the DFT-D3 method with the Becke-Johnson damping function [19]. The bulk structure was optimized with a $6\times 6\times 4$ $k$-grid within the Monkhorst–Pack scheme [47]. As a convergence criterion of the optimization loop, we took the energy change below $10^{-6}$ eV and $10^{-8}$ eV for the ionic and electronic degrees of freedom, respectively. The optimized structure was used to construct slabs with different number of PbI2 layers.
III Results & Discussion
III.1 Crystal structure
@@
The crystal structure of PbI2 is schematically depicted in Figs. 2(a) and 2(b). The PbI2 monolayer adopts a hexagonal in-plane arrangement with Pb atoms sandwiched between two layers of I atoms. Each Pb atom is surrounded by six I atoms, forming a nearly octahedral environment. The bulk structure consists of stacked layers held together by weak van der Waals forces, forming the 2H polytype, the most stable phase of PbI2 [37, 44, 16, 4, 78], which is the structure of the multilayered films studied in this work (see subsection III.3.1). We note that previous ARPES studies on bulk PbI2 have focused on the 4H polytype, in which adjacent layers are arranged in a symmetric reflection with respect to the $c$-axis [10]. Bulk PbI2 crystallizes in the $P\bar{3}m1$ space group (No. 164) with lattice parameters $a=b=4.56$ Å and $c=6.99$ Å [83]. The surface and bulk Brillouin zones (BZs) are presented in Fig. 2(c), along with their high-symmetry points.
III.2 Theoretical electronic structure
Figure 3: Orbital-resolved band structure for (a) one-, (b) two-, (c) three-, (d) seven-layer PbI2 films, and (e) the bulk compound. Grey lines represent the full DFT-calculated bands, while red and blue points indicate the projection onto the $p_{z}$ orbitals of Pb and I atoms, respectively. The band structure of the thin films refers to the 2D surface-projected Brillouin zone, while the one of the bulk compound refers to the 3D bulk Brillouin zone [see panel (c) of Fig. 2]. The origin of all energy axes has been set to the valence band maximum.
@@ III.2 Theoretical electronic structure
We will now discuss the electronic structure of PbI2 films as predicted by our DFT calculations. The most important change as a function of film thickness is the emergence of a direct band-gap, which has an immediate impact on the properties and technological applications of PbI2 [33, 64, 2]. Fig. 2(d)-(f) shows the evolution of the electronic structure with an increasing number of PbI2 layers. For a single layer [panel (a)] the band-gap is indirect due to the fact that the VBM is not located at a high-symmetry point of the surface Brillouin zone (SBZ), but rather in-between $\bar{\Gamma}$ and $\bar{\text{M}}$. As the thickness increases [Figs. 2(e) and 2(g)], the VBM shifts towards $\bar{\Gamma}$ giving rise to a direct band-gap for films of three layers (3L) and beyond. The calculated gap size slightly varies as a function of thickness, ranging from $\sim$1.95 eV for a single layer, to $\sim$1.54 eV for 7L films. Since the location of the CBM is not affected by the increasing number of layers, we underline the key role of the VBM in establishing the nature of the band-gap: as the thickness increases beyond 1L, the emergence of electronic bands with a maximum at $\bar{\Gamma}$ determines the position of the VBM, which stays at $\bar{\Gamma}$ even after the development of the bulk continuum.
@@
The orbital-resolved band-structure is shown in Fig. 3. The gray lines represent the full DFT-calculated bands, while the different colors indicate projections onto the $p_{z}$ orbitals of the Pb (red) and I (blue) atoms. As thickness increases, the bands gradually approach the bulk-like dispersion. The conduction band is dominated by Pb $p_{z}$ states, while the top of the valence band in multilayer films has an I $p_{z}$ character. In fact, as the thickness increases, interactions between the I $p_{z}$ orbitals of neighboring layers result into the emergence of the electronic bands, that constitute the VBM at $\bar{\Gamma}$ in multilayer films [80]. The shift of the VBM to $\bar{\Gamma}$ is obvious when the film thickness is at least 3L [panels (f) and (g) in Fig. 2], with the bilayer film [panel (e)] marking an intermediate stage between the monolayer behavior and the electronic structure of thicker multilayer films. The indirect-to-direct band-gap crossover of PbI2 is in agreement with previous theoretical studies [80, 71, 82, 77]. It has also been used to explain the findings of experimental measurements with no direct view on the electronic structure of PbI2  [82, 66]. However, a direct experimental verification of the indirect-to-direct band-gap crossover has never been performed and it will be the main focus of subsection III.3.
III.3 Experimental Electronic structure
III.3.1 Constant-energy surfaces
Figure 4: (a) Out-of-plane constant energy map at a binding energy of 44 eV, with green and blue arcs indicating constant photon energy contours. The thickness of the arcs approximates the $k_{z}$ broadening in ARPES and is equal to $\lambda^{-1}$, where $\lambda$ represents the photoelectron escape depth  [69]. The transformation to $k$-space was performed assuming an inner potential of $V_{0}=15$ eV. (b) and (c) Constant energy maps for monolayer and seven-layer PbI2, respectively, measured at $k_{z}=0$ (photon energy 80 eV), at a temperature of $16$ K, using linearly polarized light with horizontal polarization. Each panel presents four constant-energy contours for binding energies ranging from $5.5$ eV to $2.5$ eV, with a step of 11 eV.
The electronic structure of the 7L films being three-dimensional, one needs to identify the photon energies that allow us to probe the high-symmetry planes of the 3D Brillouin zone. To this end, we probed the constant energy contours in the plane perpendicular to the sample surface by means of photon-energy-dependent ARPES. The result is shown in Fig. 4(a) at a binding energy for which the periodicity of the experimental features was most evident. We note that the out-of-plane periodicity of the electronic structure in Fig. 4(a) ($2\pi/c=0.9$ Å-1) is in excellent agreement with the out-of-plane constant of the 2H polytype of PbI2 [4, 44, 55, 45]. Through a comparison to the theoretical electronic dispersion of bulk PbI2 along the $\Gamma-\text{A}$ direction [see Fig.  6(b) and Refs. [80, 71, 10]], we identify the high-symmetry points corresponding to incident photon energies 65 eV and 80 eV, as, respectively, the A and $\Gamma$ points of the bulk Brillouin zone. We can therefore probe the electronic structure in the high-symmetry planes $k_{z}=0$ and $k_{z}=\pi/c$ [see Fig. 2(c)] using, respectively, 80 eV and 65 eV photons. Although the electronic structure of the 1L films is, as expected, 2D (see Appendix A), when we compare it to the 7L films, we used identical photon energies in order to account for the effect of the photoemission matrix elements that are highly dependent on the experimental conditions [48].
The in-plane constant energy surfaces observed in Fig. 4(b) and  4(c) exhibit an obvious hexagonal symmetry that matches the periodicity of the surface-projected Brillouin zone (SBZ). The energy planes correspond to the same binding energy in both panels varying from $5.5$ eV (bottom) to $2.5$ eV (top). We note that, due to the insulating character of PbI2, the Fermi energy was determined by the metallic states of graphene originating from the substrate (see Appendix B). Although the overall shape of the constant-energy contours does not change substantially between the 1L and 7L films, the observed differences are due to changes of the electronic structure with increasing thickness. Such changes comprise the evolution of well-separated states into a band continuum as predicted by theory [Fig. 2(d)-(g)], slight modifications of the fine details in the energy-momentum dispersion and the relative energy shift of some electron bands (see subsection III.3.4 for more details).
III.3.2 Experimental fingerprints of the band-gap crossover
Figure 5: Electronic structure of 1 layer (1L) and 7 layers (7L) PbI2 films. (a) and (b) Band structure of, respectively, 1L and 7L films along the $\bar{\text{K}}$-$\bar{\Gamma}$-$\bar{\text{M}}$-$\bar{\text{K}}$ path. (c) and (d) Band structure of, respectively, 1L and 7L films with maximally enhanced contrast for each momentum distribution curve (MDC). The contrast enhancement was achieved by normalizing each MDC to its integrated intensity. The tail of the red arrows indicates the location of the VBM for both films. (e) and (f) Constant energy maps corresponding to a binding energy of $1.8$ eV (i.e. at the valence band maximum) for 1L and 7L films, respectively, under identical measurement conditions. (g) Energy distribution curves (EDCs) corresponding to panels (c) and (d), obtained by integrating over a momentum window of 0.1 Å-1 around $\bar{\Gamma}$. The blue (black) curve corresponds to 7L (1L) films. Measurements have been performed at 1616K using photons of $80$ eV and linear horizontal polarization.